# Large Language Models:

### v1

@@ 4.1 The Researcher’s Prediction Problem
and it recovers a particular weighted average of the LLM’s loss over all economically relevant text pieces.
With this in hand, we introduce the researcher’s inferential goal in the prediction problem. What can the researcher conclude about the LLM’s predictive performance in a research context given that the LLM is a black box?
The most relevant information the researcher has about the LLM is usually not about the different design choices made inside the black box, but rather about high-level properties regarding the model’s behavior or output. For example, builders of LLMs report their performance on natural language processing benchmarks, such as BIG-Bench or MMLU, and standardized exams, like the SAT and GRE.
We model this by assuming the researcher only knows that a given LLM satisfies some property in their research context $Q(\cdot)$; more formally, $\widehat{m}(\cdot;t)\in\mathcal{M}$, where $\mathcal{M}$ is some collection of possible text generators. We refer to the collection $\mathcal{M}$ as a guarantee, and the researcher only knows the text generator satisfies this guarantee (e.g., it could be any text generator that receives a score of at least 330 on the GRE). Given an LLM with guarantee $\mathcal{M}$, the researcher would like to draw their conclusions in the prediction problem based on the sample average loss.
@@
The large language model $\widehat{m}(\cdot;t)$ with guarantee $\mathcal{M}$ generalizes in the research context $Q(\cdot)$ if, for all text generators satisfying the guarantee $\widehat{m}(\cdot)\in\mathcal{M}$,
$\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r);Y_{r})\mid T=t\right]=\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r);Y_{r})\right].$
The large language model $\widehat{m}(\cdot;t)$ with guarantee $\mathcal{M}$ is a general-purpose technology for prediction if it generalizes in all research contexts $Q(\cdot)\in\mathcal{Q}$.
In words, the LLM $\widehat{m}(\cdot;t)$ with guarantee $\mathcal{M}$ generalizes in the research context $Q(\cdot)$ if evaluating its sample average loss recovers its average loss over the researcher’s target collection of text pieces in their research context. Notice that this is an “out-of-sample” inferential goal in the prediction problem. Importantly, under Definition 1, the researcher’s workflow in the prediction problem is justified for any LLM that satisfies the guarantee $\mathcal{M}$ — the researcher can ignore all of the details in the design of the LLM and safely proceed to draw conclusions based on the sample average loss knowing only the guarantee $\mathcal{M}$ is satisfied.
We further say that the LLM $\widehat{m}(\cdot;t)$ with guarantee $\mathcal{M}$ is a general-purpose technology for prediction if this empirical workflow is justified in all possible contexts $Q(\cdot)\in\mathcal{Q}$. Knowing the guarantee is satisfied implies this workflow in the prediction problem is justified in any research context.
To make this more concrete, let us return to our earlier empirical applications.
Example: Congressional legislation
Consider researchers that would like to predict whether a bill was passed by either house of the United States Congress $Y_{r}$ using only its short text description $r$. To do so, researchers select a given language model trained on the unknown training dataset $T=t$ and a particular prompt engineering strategy to generate responses $\widehat{m}(r;t)$. For example, researchers may prompt the language model as
Answer this question as if you were a helpful research assistant for a political scientist. Here is the description of a piece of legislation: ‘‘A bill to revise the boundary of Crater Lake National Park in the State of Oregon.’’ Will this bill pass either the United States House of Representatives or Senate? Think carefully.
Each researcher calculates the sample average loss of the LLM’s predictions on their own collected sample of Congressional bills. Different researchers may study alternative sampling distributions over Congressional bills, studying particular Congresses, for example, or particular policy areas. When can any researcher that is using a given LLM in this manner reliably conclude that whether a bill passes Congress is predictable by its description? $\blacktriangle$
Example: Financial news headlines
Consider researchers that would like to predict a company’s realized returns $Y_{r}$ based on the text of a financial news headline $r$. To do so, researchers selects an LLM trained on the unknown training dataset $T=t$ and a particular prompting strategy in order to generate predictions $\widehat{m}(r;t)$. For example, researchers may prompt the LLM as
You are a knowledgeable financial analyst. Here is a headline about Bank of New York Mellon: ‘‘Bank of New York Mellon Q1 EPS \$0.94 Misses \$0.97 Estimate, sales \$3.9B Miss \$4.01B Estimate.’’ Will this headline increase or decrease the Bank of New York Mellon’s stock price?
Each researcher calculates the sample average loss of the LLM’s predictions on their own collected sample of financial news headlines. Different researchers may select alternative sampling distributions over financial news headlines, studying particular time periods or particular industries. When can any researcher using a given LLM in this manner reliably conclude that realized returns are predictable by financial news headlines? $\blacktriangle$
4.2 Training Leakage as a Threat to Prediction
We next clarify what guarantee $\mathcal{M}$ is necessary and sufficient for an LLM to generalize in a research context, and therefore whether it is a general-purpose technology for prediction.

### v3

@@ 3.1 The Researcher’s Prediction Problem
$\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r),Y_{r})\mid T=t\right]=\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r),Y_{r})\right].$
This definition formalizes an out-of-sample prediction goal. The LLM generalizes if its sample average loss reflects its predictive performance on the target population. Under Definition 1, the researcher’s workflow in the prediction problem is justified for any LLM satisfying the guarantee $\mathcal{M}$. The researcher can draw conclusions based on the sample average loss knowing only the guarantee $\mathcal{M}$ is satisfied.
Example: Congressional legislation
Consider researchers predicting whether a bill passes either house of Congress $Y_{r}$ using only its text description $r$. The LLM generates predictions $\widehat{m}(r;t)$ using a specific prompting strategy. Each researcher calculates the sample average loss of the LLM’s predictions on their own collected sample of Congressional bills. Different researchers face different prediction problems, formalized by alternative research contexts $Q(\cdot)$. One researcher might assess whether the LLM can predict outcomes for future legislation by sampling bills after a cutoff date, such as all bills from 2025 onward. Another might evaluate whether the LLM’s predictions generalize to novel policy domains, sampling bills on as cryptocurrency regulation or artificial intelligence policy. $\blacktriangle$
Example: Financial news headlines
Consider researchers predicting a company’s realized returns $Y_{r}$ from a financial news headline $r$. The LLM generates predictions $\widehat{m}(r;t)$ using a specific prompting strategy. Each researcher calculates the sample average loss on their collected headlines. Different researchers again face different out-of-sample prediction problems, formalized by alternative research contexts $Q(\cdot)$. One researcher might assess whether the LLM predicts returns during future market conditions, sampling headlines from 2025 onward. Another might evaluate whether predictions generalize across company types, focusing on small-cap technology firms or international equities. $\blacktriangle$
3.2 Training Leakage as a Threat to Prediction
Using LLMs in prediction problems requires one condition: no training leakage between the LLM’s training data and the researcher’s evaluation sample.
Lemma 1.
[…]
$\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))\right]=\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))\mid T=t\right]-\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\left(\frac{q_{r}^{T\mid D}(t_{r})}{q^{T}_{r}(t_{r})}-1\right)\ell(Y_{r},\widehat{m}(r))\right].$
Proposition 1.
The LLM $\widehat{m}(\cdot;t)$ generalizes for research context $Q(\cdot)\in\mathcal{Q}$ if and only if it satisfies the guarantee $\mathcal{M}(Q)$ for
$\mathcal{M}(Q)=\left\{\widehat{m}(\cdot)\colon-\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\left(\frac{q^{T\mid D}_{r}(t_{r})}{q^{T}_{r}(t_{r})}-1\right)\ell(Y_{r},\widehat{m}(r))\right]=0\right\}.$ (4)
The no-training leakage condition (Equation 4) captures the extent to which overlap between the LLM’s training data and the researcher’s sample covaries with the LLM’s predictive performance. It has an intuitive interpretation as omitted variable bias. The term $\frac{q^{T\mid D}_{r}(t_{r})}{q^{T}_{r}(t_{r})}-1$ measures how learning that the researcher sampled text piece $r$ updates beliefs about whether $r$ appeared in the LLM’s training data. When this correlation is positive – text pieces in the researcher’s sample are more likely to have been in the training data – and the LLM performs well on such pieces, the overall bias term is positive and the sample average loss overstates true predictive performance. Uncertainty over what entered into the LLM’s training data acts like an omitted variables bias in the prediction problem.
This bias arises from both the researcher’s choice of out-of-sample prediction and our beliefs about what entered into the LLM’s training data. Section 3.3 discusses how researchers can prevent training leakage through careful choices of model and prediction exercise.
3.2.1 Some Evidence of Training Leakage
Training leakage is a well-documented problem in computer science. LLMs’ training datasets frequently contain examples from popular benchmark evaluations (sainz2023nlpevaluationtroubleneed; golchin2024timetravelllmstracing; lm_index). This has generated skepticism about evaluating LLMs on any publicly available data (e.g., ravaut2024largelanguagemodelscontaminated). But this evidence focuses on CS applications. Might economic prediction problems face similar risks?
Assessing Training Leakage in Congressional Legislation
We assess training leakage in an empirical setting relevant for economists studying politics and political economy: congressional legislation. The Congressional Bills Project (CAP; CBP) contains descriptions $r$ for over 400,000 bills proposed in Congress, along with whether each passed the House or Senate $Y_{r}$. We sample $10,000$ bills introduced from 1973 to 2016 and test whether passage can be predicted from text descriptions alone — a potentially challenging task given the strategic dynamics of congressional voting. Among these bills, only 7.4% pass the House and 6.0% pass the Senate.
We generate predictions $\widehat{Y}_{r}=\widehat{m}(r;t)$ based on each bill’s description $r$ by prompting GPT-4o (see Appendix Figure A13 for the specific prompt). GPT-4o correctly predicts the bill’s outcome 91.2% of the time in the House and 92.5% of the time in the Senate (left panel of Appendix Table 1(b)). What drives the model’s ability to accurately predict whether a Congressional bill will pass the House or the Senate based only on its description? The answer is that the text of congressional legislation is likely included in its training dataset.
To evaluate training leakage, we prompt the model to complete each bill’s text description based on only the first half of its text, following research in computer science such as golchin2024timetravelllmstracing (see Appendix Figure A14 for the specific prompt). If a model reproduces the text exactly, it has likely seen it during training. This is not a necessary condition for training leakage; models may benefit from exposure to a text piece without memorization. Perfect reproduction nonetheless offers compelling evidence.
On 344 bills, GPT-4o completes the bill’s text description exactly as it is written, indicating that not only was GPT-4o likely trained on these text pieces but it appears to have memorized them. Appendix Figure A1 provides two examples of successful completions. On the other bills, GPT-4o’s completed bill descriptions are close to the original bill descriptions. Word embeddings of GPT-4o’s descriptions are far closer to those of the originals than the average distance between the word embeddings of two randomly selected bills (left panel of Appendix Table 2(b)).
One might wonder whether this training leakage could be addressed through prompt engineering — for example, by commanding the models to not pay attention to any information past a certain date. Applying this prompt engineering strategy to our sample of Congressional bills (see Appendix Figure A13 and A14 for associated prompts), we still find substantial evidence of training leakage. Even when explicitly told to not consider any information past the bill’s introduction date in Congress, GPT-4o can still accurately predict its outcome in the House and the Senate based on these small snippets of text (right panel of Appendix Table 1(b)). GPT-4o still exactly completes nearly the same number of bill descriptions as without the prompt engineering (330 versus 344), and the word embeddings of its completed descriptions remain quite close on average to those of the originals (right panel of Appendix Table 2(b)).
Assessing Training Leakage in Financial News Headlines
We consider another domain relevant for economists: financial markets. Existing work (e.g., glasserman2023assessinglookaheadbiasstock; lopezlira2024chatgptforecaststockprice) found that LLMs predict stock returns accurately from news headlines. We test whether this reflects training leakage using publicly available headline data (Headlines) covering nearly 4 million headlines for 6,000 publicly traded companies from 2009-2020.
We sampled 10,000 financial news headlines from 2019, and we prompt GPT-4o to complete each financial news headline based on only 50% of its text (see Appendix Figure A15 for the specific prompt). GPT-4o reproduces 60 financial news headlines exactly as they were written in the publicly available dataset, indicating that GPT-4o was likely trained on these headlines and memorized them. Appendix Figure A2 provides two examples of successfully completions. On all other headlines, word embeddings of the model’s completions are close to those of the original headlines (left panel of Appendix Table 3(b)).
We again explore whether explicitly incorporating date restrictions into the LLM prompt moderate this evidence of training leakage (see Appendix Figure A15 for the associated prompts). [^6] See also wongchamcharoen2025largelanguagemodelsllms. Another prompting strategy – entity masking – aims to prevent memorized information by masking identifiers in prompts (glasserman2023assessinglookaheadbiasstock; sarkar2024lookahead; EnglebergEtAl(25)-EntityNeutering). However, implementation details matter, and its wider applicability is unclear. In our congressional legislation example, it is not obvious what should be masked. Surprisingly, this appears to make the problem worse; GPT-4o now reproduces 73 headlines exactly, and word embeddings of GPT-4o’s completions with the date restriction are still on average close to those of the original headlines.
sarkar2024lookahead provide additional evidence of training leakage in finance: prompting Llama 2 to predict risks from September-November 2019 earnings calls, the LLM mentions Covid-19 in over 25% of cases—a form of ”look-ahead bias” from training on future information. Combined with our findings and other evidence (glasserman2023assessinglookaheadbiasstock; lopezlira2025memorizationproblemtrustllms), this indicates substantial risk of training leakage in finance applications.
3.3 Practical Guidance for Prediction Problems
The no-training-leakage condition (Equation 4) may seem abstract, but it provides concrete guidance for empirical practice. Researchers must consider what population they are predicting on, how their evaluation sample relates to it, and how their evaluation sample relates to the LLM’s training corpus.
To make this concrete, recall that the bias term in Proposition 1 depends on $\frac{q^{T\mid D}(t_{r})}{q^{T}(t_{r})}-1$ – how much does learning that the researcher sampled text piece $r$ update our beliefs about whether $r$ was in the training data. No training leakage requires this term to be zero, or at least uncorrelated with the LLM’s performance. Different prediction problems and model choices achieve this in different ways. We illustrate this by considering the threat of training leakage and how to manage it across several examples.
Example: Lookahead Bias and Time-Stamped Models
Consider a researcher who wants to predict stock returns from future financial news headlines. If the researcher evaluates on recent headlines from 2024 using an LLM trained on data through 2025, the leakage term is almost certainly positive: both the researcher and the LLM builders had access to 2024 headlines. Any predictive success may reflect training leakage.
The solution in this case is to use open-source LLMs with fixed published weights, such as the Llama family (touvron2023llama2openfoundation; dubey2024llama3herdmodels) and others (e.g., BLOOM, OLMo, etc.), or time-stamped training data (e.g., Sarkar(24); he2025chronologicallyconsistentlargelanguage). If the researcher uses a model with weights published on date $\tau$ and constructs an evaluation sample using documents published after $\tau$, then $q^{T}_{r}(t_{r})=0$ mechanically since they could not have been in the training data. This eliminates training leakage by design.
Closed models like the GPT family from OpenAI or the Claude family from Anthropic do not permit this solution. Their training data is undisclosed, and they may be continuously fine-tuned. Researchers in natural language processing now regularly caution against sending test data to the APIs or chat interfaces of closed LLMs since these data may be used in further fine-tuning or the development of new models (jacovi-etal-2023-stop).777balloccu2024leakcheatrepeatdata estimates that 263 benchmarks may have been inadvertently leaked to OpenAI through use of the chat interface and API, and cheng2024dateddatatracingknowledge questions the validity of publicly stated “knowledge cutoffs” in closed LLMs. Most worryingly, BarrieEtAl(24)-LLMReplication illustrates that the results of submitting the same prompt to GPT-4o yields results that change month-to-month, despite there being no publicly announced changes to the underlying model. $\blacktriangle$
Example: Prediction on Confidential Documents
Suppose a researcher wants to predict, for example, whether administrative case notes are predictive of some decision, like pretrial release. In examples like this, the target population consists of confidential documents that were never publicly released. Since these documents never entered any public corpus, we can again credibly argue that $q^{T}_{r}(t_{r})=0$ for all documents since they were never available for training. This reasoning hinges on credibly claiming the documents’ provenance. $\blacktriangle$
Example: Random Sampling from a Known Corpus
Consider a researcher with a complete corpus of economics papers published between 2000 and 2020 who wants to predict citation counts from abstracts. She draws a random sample to form her evaluation set. In this case, the researcher’s sampling process is independent of the LLM’s training data: her random number generator knows nothing about which papers OpenAI included in their training data. This implies $q^{T\mid D}_{r}(t_{r})=q^{T}_{r}(t_{r})$ – learning that a paper was randomly selected provides no information about whether it was trained on — so the leakage term equals zero.
The logic is familiar from causal inference: randomization eliminates omitted variable bias by making treatment assignment independent of potential outcomes. Here, random sampling eliminates training leakage by making the researcher’s evaluation sample independent of the LLM’s training data. This can be applied whenever the researcher draws a genuinely random sample from a well-defined corpus before using an LLM. The converse is also important: just as non-random treatment assignment resurrects possible confounding in causal inference, non-random sampling may resurrect training leakage. $\blacktriangle$
Example: Prediction on the Complete Population
As a final example, suppose the researcher collected all Congressional bill descriptions from 1973–2016 and wants to understand what textual features predict passage. The researcher seeks to descriptively characterize patterns in this specific historical corpus, not to make predictions on some population of legislation. When the researcher observes the entire population, the prediction problem changes character. There is no sampling, and hence no inference to a broader population. In this case, $q^{D}_{r}=1$ and the leakage condition is satisfied trivially: we are not asking whether performance generalizes beyond what we observe. The complete-population case licenses only descriptive conclusions about the collected corpus. $\blacktriangle$
More broadly, these examples illustrate a general principle: no training leakage is ultimately about both research design and model properties. Researchers must be explicit about their target population, their sampling procedure, and how that procedure relates to possible training corpora. When in doubt, the safest approach combines open-source models with published weights or documented training data and evaluation samples constructed to mechanically exclude the model’s training corpus.
4 Estimation with Large Language Models
In estimation problems, the researcher measures economic concepts $V_{r}$ expressed in text pieces $r$ to estimate downstream parameters. The measurement $f^{*}(\cdot)$ is costly to scale across thousands or millions of text pieces. This is where LLMs enter as potential substitutes.
Using LLM outputs in plug-in estimation requires a strong assumption: the LLM must reproduce the existing measurement everywhere. As Section 2.3 discussed, LLM performance varies across similar tasks, and benchmark evaluations provide little guidance on new applications. We demonstrate this empirically: seemingly minor choices – which LLM, which prompt – substantially affect downstream parameter estimates in applications to finance and political economy. These choices change the magnitude, significance, and even sign of estimated parameters. The solution: collect a small validation sample and use it to debias the LLM’s outputs.
4.1 The Researcher’s Estimation Problem
The researcher specifies a parameter $\theta\in\Theta$ and a moment condition $g(\cdot)$ that identifies this parameter. If she collected the economic concept $V_{r}$ on each text piece $r$, she would calculate
For example, letting $g(V_{r},W_{r};\theta)=(V_{r}-W_{r}^{\prime}\theta)^{2}$, the researcher studies how $V_{r}$ relates to the linked variables $W_{r}$, and $\widehat{\theta}^{*}$ is the sample regression coefficient. Due to the text processing problem, however, the researcher cannot calculate $\widehat{\theta}^{*}$ directly.
The researcher prompts an LLM to measure the economic concept on each text piece, $\widehat{V}_{r}:=\widehat{m}(r;t)$, and plugs in the LLM’s labels
When can we draw conclusions about $\widehat{\theta}^{*}$ based on $\widehat{\theta}$?
We associate the researcher with a research context $Q(\cdot)\in\mathcal{Q}$ satisfying Assumption 1. We assume that the researcher could study any moment condition $g(\cdot)\in\mathcal{G}$ satisfying the following assumption.
Assumption 2.
For all $g(\cdot)\in\mathcal{G}$, $g(\cdot)$ is differentiable and there exists some $\overline{G}>0$ such that $\left|\frac{\partial g(v,W_{r};\theta)}{\partial\theta}\right|\leq\overline{G}$ for all $r\in\mathcal{R}$, $\theta\in\Theta$, and values of the economic concept $v$.
As the number of economically relevant text pieces grows large (see Appendix C.1), we can characterize the behavior of both estimators. The target moment condition converges to, at any parameter value $\theta\in\Theta$,
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}g(V_{r},W_{r};\theta)-\frac{1}{\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}]}\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}g(V_{r},W_{r};\theta)\right]\xrightarrow{p}0.$ (7)
Conditional on the LLM’s realized training dataset, the plug-in moment condition converges to, at any parameter value $\theta\in\Theta$,
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}g(V_{r},W_{r};\theta)-\frac{1}{\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\mid T=t]}\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}g(\widehat{m}(r;t),W_{r};\theta)\mid T=t\right]\xrightarrow{p}0.$ (8)
As in the prediction problem, conditioning on the training dataset simplifies analysis by treating the LLM as a fixed mapping.
Given an LLM with guarantee $\mathcal{M}$, the researcher would like to recover the moment condition defined using the economic concept.
@@ 4.1 The Researcher’s Estimation Problem
[…]
The LLM $\widehat{m}(\cdot;t)$ with guarantee $\mathcal{M}$ is a general-purpose technology for estimation if it automates the researcher’s measurement process for all $g(\cdot)\in\mathcal{G}$ and $Q(\cdot)\in\mathcal{Q}$.
The excitement around LLMs stems from their potential as “general-purpose technologies” – tools deployable across diverse applications without task-specific engineering (e.g., ElondouEtAl(24)). For estimation problems, this would mean reliably substituting for existing measurements across different economic concepts and research contexts. Definition 2 formalizes what this requires: the LLM must produce moment conditions matching the existing measurement procedure, regardless of which moment condition or population the researcher studies. If an LLM satisfies this property, researchers could confidently use its outputs for plug-in estimation knowing only the guarantee $\mathcal{M}$.
4.2 Measurement Error as a Threat to Estimation
We clarify what guarantee $\mathcal{M}$ is necessary and sufficient for an LLM to automate the existing measurement. We decompose the difference between the plug-in moment condition and the target moment condition into two terms.